High-power experimental engines are being developed by the Stevens Motor Company for use in its new sports coupe. The engineers have calculated the maximum horsepower for the engine to be 750HP. Twenty five engines are randomly selected for horsepower testing. The sample has an average maximum HP of 770 with a standard deviation of 75HP. Assume the population is normally distributed.

Step 1 of 2 :  

Calculate a confidence interval for the average maximum HP for the experimental engine. Use a significance level of α=0.1α=0.1. Round your answers to two decimal places.
Answer #1

df = n - 1 = 25 - 1 = 24

t critical value at 0.1 significance level with 24 df = 1.711

90% confidence interval for is

- t * S / sqrt(n) < < + t * S / sqrt(n)

770 - 1.711 * 75 / sqrt(25) < < 770 + 1.711 * 75 / sqrt(25)

744.34 < < 795.67

90% CI is ( 744.34 , 795.67 )